OpenStudy (luvhawks711):

3. A. Graph: B. Is this a growth or decay function? C. State the percent of growth or decay. 4. A. Graph: B. Is this a growth or decay function? C. State the percent of growth or decay.

OpenStudy (luvhawks711):

3, y=3(2)^x 4.f(x)=10(0.4)^x

OpenStudy (luvhawks711):

@Kyliejones @freemap